OpenStudy (anonymous):

Identify the correct slope and y intercept of the equation 6x - 3y = 12. slope = 2; y intercept at (0, -4) slope = -2; y intercept at (0, 4) slope = ; y intercept at (0, 15) slope = ; y intercept at (0, 9) @mulline ?

OpenStudy (anonymous):

@hartnn @radar ?

hartnn (hartnn):

try to isolate 'y' from that equation. bring it in the form of y=mx+c

hartnn (hartnn):

then m= slope and c = y-intercept

OpenStudy (anonymous):

I tried and I got the completely wrong answer.

OpenStudy (anonymous):

I got like.. y= x + 4

hartnn (hartnn):

how? try again ?

OpenStudy (anonymous):

I don't even know. But ok, I'll try again.

hartnn (hartnn):

6x - 3y = 12

hartnn (hartnn):

add 3y on both sides , what u get ?

OpenStudy (anonymous):

Ohhhh, I accidentally subtracted.

OpenStudy (anonymous):

Thanks, I think I got it. :)

hartnn (hartnn):

welcome ^_^